# SAQ-SAM: Semantically-Aligned Quantization for Segment Anything Model

### v1

@@ 3.3 Prompt-Aware Reconstruction
where $O^{l},{\hat{O}}^{l}$ are the outputs of $l$-th block from the FP model and quantized model, respectively. $\alpha$ is the adaptive rounding factor, an extra weight quantization parameter introduced by AdaRound [26]. While this approach effectively enhances PTQ performance on conventional models, it struggles with the SAM model, particularly due to severe overfitting to pure image information. In the SAM model, to achieve the desired result, the image embedding produced by the image encoder is utilized to interact with prompt tokens in the mask decoder. However, simply considering local visual response overlooks the prompt’s intent, potentially adding redundant information that disrupts the interaction between the prompt and image features, resulting in distorted segmentation results.
Interaction Response Reconstruction. To address the aforementioned issue, we incorporate the interaction between prompt and image tokens into the reconstruction process, enabling more effective matching under the supervision of the FP model. Specifically, instead of aligning the local visual response, we reconstruct the hybrid image tokens that incorporate the prompt information by leveraging the interaction in the mask decoder.
As illustrated in Figure 2, we utilize the off-the-shelf interaction module in SAM’s mask decoder to incorporate prompt information for the image tokens, without additional components or extra training. In this way, the hybrid image tokens can be obtained as follows:
$T_{ip}^{k}=\text{MaskDecoderModule}(T_{i}^{k},T_{p}),$ (10)
where $T_{p}$ denotes prompt tokens encoded by prompt encoder, $T_{i}^{k}$ denotes image tokens derived from the output of the $k$-th stage. MaskDecoderModule refers to Two-Way Transformer, which is the primary interaction module in the mask decoder. Subsequently, these tokens are reconstructed to learn the quantization parameters, i.e., minimizing the L2 distance between the hybrid image tokens and the corresponding response in the FP model as follows:
@@
Note that to ensure efficiency of the prompt-image interactions, the defined image tokens $T_{i}^{k}$ skip subsequent layers and directly perform dimensionality reduction through the neck. This approach avoids prohibitively high computational cost associated with complete forward passes and gradient backpropagation, and thus only introduces a slight additional overhead compared to traditional local response reconstruction. Moreover, in addition to improving efficiency, the layer-skipping strategy also offers advantages in performance retention. By shortening optimization paths, it better preserves local distributional properties and reduces the loss of interaction information that may occur due to excessively long propagation processes.
To validate the effectiveness of utilizing these skip-layer tokens for interaction, we conduct verification experiments in which these tokens are used as the final visual embeddings to participate in the mask decoder inference. The segmentation results, as shown in Figure 4, demonstrate that the outputs from each stage are capable of producing rational segmentation results without further processing by subsequent layers. This provides implicit evidence that the proposed approach enables an effective interaction, with the outputs of each stage functioning as image embeddings at different levels of granularity.
@@
To further enhance the efficiency of parameter learning and improve overall performance, we partition the Transformer layers into multiple stages and adopt a stage-wise learning approach for quantization parameters, where the parameters within each stage are jointly optimized. Previous work [29] has demonstrated the superiority of jointly optimizing multiple quantization blocks in reconstruction, as it effectively captures weight correlations across blocks. Building on this insight, we intuitively define global attention layers as boundaries for stage partitioning. For example, in SAM-B, layers $L2,L5,L8$ and $L11$ employ global attention, while the remaining layers utilize window attention. Accordingly, layers $\{L0,L1,L2\}$ constitute stage 0, and this pattern repeats for subsequent stages.
4 Experiment
@@ 4.1 Experimental Setup
Tasks and Datasets. Our experiments focus on two prevalent SAM practices: instance segmentation and semantic segmentation. In instance segmentation tasks, SAM generates segmentation masks according to provided bounding boxes. The manual prompting process is substituted by a detector that identifies the target objects within the image and produces the corresponding bounding boxes’ coordinates for SAM. We evaluate the performance of the quantized SAM on the MS-COCO [19] dataset, with the mean Average Precision (mAP) serving as evaluation metric. In the semantic segmentation experiment, SAM is employed to enhance the quality of segmentation masks produced by traditional semantic segmentor. A large number of points, which are evenly distributed across the image, act as prompts. These prompts are fed into SAM in multiple bathches, generating a plenty of masks. Through a series of post-processing steps, these masks are matched with the outputs of semantic segmentor, thus forming the final panoptic segmentation result [1]. The evaluation dataset is ADE20K [40], with mean Intersection over Union (mIoU) serving as performance metric.
@@
Implementation Details. We take PTQ4SAM [23] as the baseline method, while maintaining largely consistent experimental settings for fair comparison. To be specific, the quantization scheme employs per-tensor asymmetric quantization for activations and per-channel asymmetric quantization for weights, with exceptions for the first and last layers which remain unquantized. The calibration set contains 32 images randomly sampled from the training dataset. For the Perception-Consistency Clipping (PCC) method, we set threshold factor $\theta$ to 0.5. All QK activations (i.e., the inputs and outputs of Query Linear and Key Linear) are calibrated based on Attention Focus Overlap metric using the first sample, while other activations are calibrated based on Mean Squared Error (MSE) using all samples. The weights are calibrated using the MSE. For Prompt-Aware Reconstruction (PAR), the image encoder employs stage-wise learning, while the mask decoder employs layer-wise learning with 2000 iterations. Exceptionally, the additional cross-attention block that is equipped at the end of the mask decoder learns 10000 iterations due to its large loss. In general, compared to PTQ4SAM settings, which carry block-wise learning with 20000 iterations, our method requires much lower time cost.
Baseline Methods. We use the advanced PTQ4SAM as our baseline [23]. For comparative analysis, we categorize the methods into two groups according to whether there is a learning process or not. For the first group of statistic-based PTQ methods, we integrated the PCC method into PTQ4SAM-S [23], and this is refered to as SAQ-PCC. As for the second group of learning-based PTQ methods, We build on PTQ4SAM-L [23] and substitute the QDrop [32] with our PAR method, which is abbreviated as SAQ-PAR. When the PCC method is futher incorporated as a pre-processing step, our complete method, denoted as SAQ-SAM, is established.
[…]
SAM-B SAM-L SAM-H
Detector Methods REC FP W6A6 W4A4 FP W6A6 W4A4 FP W6A6 W4A4
PTQ4SAM-S [23] 15.4 - 35.7 18.1 36.0 24.1
SAQ-PCC 29.8 2.8 35.7 20.9 36.1 25
PTQ4SAM-L [23] 30.3 16.0 35.8 28.7 36.5 33.5
SAQ-PAR 31.9 23.2 35.8 34.6 36.7 35.9
Faster R-CNN [28]
SAQ-SAM
(PCC+PAR)
33.4 32.0 27.7 36.4 35.9 34.5 37.2 36.6 35.9
PTQ4SAM-S [23] 17.4 - 40.0 20.6 40.3 26.7
SAQ-PCC 34.0 2.1 40.0 24.2 40.2 28.4
PTQ4SAM-L [23] 34.3 18.4 40.3 31.6 40.7 37.6
SAQ-PAR 35.7 26.2 40.2 38.9 40.8 39.4
YOLOX [6]
SAQ-SAM
(PCC+PAR)
37.0 35.9 30.3 40.4 40.3 39.0 41.0 40.9 39.9
PTQ4SAM-S [23] 17.9 - 41.0 20.9 41.3 27.3
SAQ-PCC 36.4 2.6 41.1 24.3 41.3 28.2
PTQ4SAM-L [23] 35.1 17.3 41.2 32.1 41.6 38.4
SAQ-PAR 37.0 29.4 41.2 39.6 41.9 40.8
H-Deformable-DETR [9]
SAQ-SAM
(PCC+PAR)
38.2 37.1 31.8 41.5 41.3 39.8 42.0 41.8 40.7
PTQ4SAM-S [23] 20.4 - 47.7 23.1 48.1 30.5
SAQ-PCC 39.4 3.5 48.0 27.8 48.2 31.6
PTQ4SAM-L [23] 40.4 14.4 48.3 36.6 48.7 43.9
SAQ-PAR 42.3 30.2 48.3 46.1 48.8 47.4
DINO [37]
SAQ-SAM
(PCC+PAR)
44.5 42.4 33.8 48.6 48.3 46.3 49.1 48.9 47.4
4.2 Instance Segmentation Results
@@
In the task of instance segmentation, we provide a detailed comparison of the performance on various sizes of SAM, including SAM-B, SAM-L and SAM-H, when used in conjunction with different detectors (including Faster R-CNN [28], YOLOX [6], H-Deformable-DETR [9], and DINO [37]). The results on COCO dataset are listed in Table 1. Our method consistently shows superior performance. Compared to baseline PTQ4SAM-S [23], SAQ-PCC achieves a nearly twofold improvement in mAP for 6-bit SAM-B (e.g., from 20.4% to 39.4% with DINO) and recovers the collapsed accuracy to an average of 2.7% mAP at 4-bit. For SAM-L and SAM-H, the advantages are particularly evident in low-bit settings. For example, SAQ-PCC increases the mAP of 4-bit SAM-L from 23.1% to 27.8% with DINO, and increases the mAP of 4-bit SAM-H from 26.7% to 28.4% with YOLOX. In terms of reconstruction methods, compared to PTQ4SAM-L [23] that uses naive QDrop, our methods achieve significant improvement with lower computational costs. For example, SAQ-PAR improves the 6-bit SAM-B by 1.6% mAP (from 30.3% to 31.9%), and the 4-bit SAM-B by 7.2% (from 16.0% to 23.2%), both with Faster R-CNN. For larger models, it boosts 4-bit SAM-L by 9.5% (from 36.6% to 46.1%) with DINO. Furthermore, if used as a pre-processing step, PCC can provide more favorable initial values for PAR optimization, thereby further enhancing the performance of low-bit quantization. For example, SAQ-SAM improves the 4-bit SAM-B by 14.5% mAP (from 17.3% to 31.8% with H-Deformable-DETR) and achieves near-lossless accuracy for 4-bit SAM-L and SAM-H(e.g., with only a drop of 1.4% for SAM-L with YOLOX, and 1.3% for SAM-H with Faster R-CNN).
Table 2: Quantization results of semantic segmentation on ADE20K datasets with SegFormer providing category labels.
Model Prec. Method REC mIoU(%)
FP - - 33.17
W6A6 PTQ4SAM-S [23] 31.16
W6A6 SAQ-PCC 32.90
W6A6 PTQ4SAM-L [23] 32.65
W6A6 SAQ-SAM 33.04
W4A4 PTQ4SAM-S [23] 31.08
W4A4 SAQ-PCC 31.29
W4A4 PTQ4SAM-L [23] 31.85
SAM-B W4A4 SAQ-SAM 32.53
4.3 Semantic Segmentation Results
In the task of semantic segmentation, we employed SegFormer [34] as the semantic branch, which provides category labels for SAM’s masks. The test results on the ADE20K dataset are reported in Table 2. Since the masks generated by SAM are exposed to a series of post-processing procedures, the harm caused by quantization on the final performance is not as obvious. Nevertheless, our methods improve the accuracy of quantized SAM-B comprehensively. For example, SAQ-PCC increases the accuracy of 6-bit quantization by 1.74% mIoU (from 31.16% to 32.9%), even surpassing the learning-based PTQ4SAM-L (with 32.65% mIoU). Furthermore, SAQ-SAM boosts the performance close to FP, with only a loss of 0.13% mIoU at 6 bits and a loss of 0.64% at 4-bit.
@@ 4.4 Ablation Study
4.4.1 Attention Focusing Threshold Setting
In the proposed PCC module, we use the threshold factor $\theta$ to define the focus area of attention. For generality, we typically set $\theta$ to 0.5. To explore the impact of this hyperparameter on the performance of our method, we carry out comparative experiments on the instance segmentation task. Setting $\theta$ to values ranging from 0.3 to 0.9, the results on 6-bit SAM-B and 4-bit SAM-L are shown in Figure 5. Our analysis indicates two key points: First, with changes in $\theta$, the performance of our clipping method slightly fluctuates, but it consistently outperforms the baseline method, with 0.5 emerging as a modest choice. Second, our method demonstrates exceptional robustness to the setting of $\theta$, achieving satisfactory performance across a range of values from 0.4 to 0.7.
Figure 5: Curve plots of ablation experiment results for attention focus threshold $\theta$. The performance of our method is robust to this hyperparameter, and the result consistently outperform the baseline, with 0.5 being a modest choice.
4.4.2 Comparison of Reconstruction Method
Inspired by  [29], we partition several layers into a stage based on the global attention layer, and treat these stages as optimization units. This scheme not only reduces the computation cost, but also captures the dependencies among blocks, which achieves superior performance. To investigate the effectiveness of our PAR method under different reconstruction granularities (per-layer and per-stage), we conduct detailed ablation experiments, comparing the performance of using QDrop reconstruction, our Prompt-Aware Reconstruction, and the cases where PCC is used as prior processing. We quantized SAM-B for instance segmentation with YOLOX as the detector. The experimental results are shown in the Table 3. Note that, for fair comparison, the learning settings across all experiments are kept consistent, including learning iterations, learning rate, optimizer and drop probability. The results show that compared with per-layer learning, consistently exhibits advantages across various methods, indicating that the per-stage granularity is a more effective choice. Furthermore, our method demonstrates superior performance under both reconstruction granularities, highlighting the contribution of introducing interactions into the reconstruction process.
Table 3: Ablation study for reconstruction methods using different optimization unit partitions. The per-stage reconstruction exhibits a greater superiority and our method outperforms the baseline at both reconstruction granularities
Prec. Method Per-Layer Per-Stage
QDrop 8.1 17.9
PAR 24.4 26.2
QDrop+PCC 16.5 26.6
W4A4 PAR+PCC 28.6 30.3
QDrop 33.7 35.6
PAR 35.7 35.7
QDrop+PCC 34.4 35.8
W6A6 PAR+PCC 35.8 35.9

### v2

@@ 3.3 Prompt-Aware Reconstruction
where $O^{l},{\hat{O}}^{l}$ are the outputs of $l$-th block from the FP model and quantized model, respectively. $\alpha$ is the adaptive rounding factor, an extra weight quantization parameter introduced by AdaRound (adaround). While this method effectively enhances PTQ performance on conventional models, it struggles with the SAM model due to severe overfitting to pure image information. Specifically, Designed to follow the user’s prompt, SAM’s image embeddings interacting with prompts embeddings in mask decoder. Therefore, simply considering local visual response overlooks the prompt’s intent, potentially introducing redundant information that disrupts the image-prompt interaction, resulting in distorted segmentation results.
Interaction Response Reconstruction. To address the aforementioned issue, we incorporate the interaction between prompt and image tokens into the reconstruction process, enabling effective image-prompt matching under the supervision of the FP model. Specifically, instead of aligning the raw visual response, we reconstruct the hybrid image tokens that integrate the prompt information by interacting in the cross-attention of mask decoder.
As illustrated in Figure 4, we utilize the off-the-shelf cross attention module in SAM’s mask decoder to incorporate prompt information for the image tokens, without additional components or extra training. In this way, the hybrid image tokens can be obtained as follows:
$T_{ip}^{k}=\text{TwoWayTransformer}(T_{i}^{k},T_{p}),$ (10)
where $T_{p}$ denotes prompt tokens encoded by prompt encoder, $T_{i}^{k}$ denotes image tokens derived from the outputs of the $k$-th stage. TwoWayTransformer refers to the Two-Way Transformer module detached from the mask decoder, which is designed for efficient image-prompt interaction. Subsequently, these tokens are reconstructed to learn the quantization parameters, i.e., minimizing the L2 distance between the hybrid image tokens and the corresponding FP response as follows:
@@
Layer Skipping Interaction. To boost the efficiency of reconstruction, we customize two designs based on SAM’s characteristics. First, we partition the Transformer layers into multiple stages and adopt stage-wise learning, where the quantization parameters within each stage are jointly optimized. Previous work (interactions-across-block) has demonstrated the superiority of jointly optimizing multiple quantization blocks in reconstruction, as it effectively captures weight correlations across blocks. Building on this insight and SAM’s architecture, we define global attention layers as boundaries for stage partitioning. For example, in SAM-B, layers $L2,L5,L8$ and $L11$ employ global attention, while the remaining layers utilize window attention. Accordingly, layers $\{L0,L1,L2\}$ constitute stage 0, and this pattern repeats for subsequent stages.
Furthermore, to promote efficient image-prompt interactions, the outputs of learning stage skip subsequent layers and directly pass the neck for interaction:
$T_{i}^{k}=\text{Neck}((\textstyle\prod_{i=0}^{k}\text{Stage}^{k})(E_{i})),$ (12)
where $E_{i}$ denotes patch embeddings, $\text{Stage}^{k}$ denotes the k-th stage, $\prod_{i=0}^{k}\text{Stage}^{k}$ represents the sequential composition of stages from stage 0 to stage k, Neck denotes SAM’s Neck module for dimension reduction of image features. This layer-skipping strategy avoids prohibitively high computational cost of complete forward passes and gradient backpropagation, thus introducing slight additional overhead compared to traditional local response reconstruction. In addition to efficiency benefits, skipping deeper layers also brings performance advantages. With shortened optimization paths, it not only preserves local distribution properties, but also avoids the potential loss of interactive information due to long propagation processes.
To validate the feasibility of using these layer-skipping tokens directly for interaction, we tentatively use them as the final visual representation for mask prediction. The results shown in Figure 5 indicate that these immature tokens can produce rational segmentation results without further processing by deeper layers. This provides implicit evidence that the layer-skipping design is capable of effective interaction, with the outputs of each stage functioning as image embeddings at different levels of semantic granularity.
Table 1: Quantization results of instance segmentation on COCO dataset with various detectors providing prompts. We abbreviate ‘Precision’ as ‘Prec.’, ‘W/A’=$x/y$ denotes the weights and activations are quantized to $x$ bit and $y$ bit, respectively, and ‘-’ indicates the mAP is below 1. Rows with green background indicate statistic-based methods, while blue background indicate learning-based methods. The bolded values represent the best-performing results under the same settings.
Detector Faster R-CNN YOLOX H-Deformable-DETR DINO
Model SAM-B SAM-L SAM-H SAM-B SAM-L SAM-H SAM-B SAM-L SAM-H SAM-B SAM-L SAM-H
FP 33.4 36.4 37.2 37.0 40.0 41.0 38.2 41.5 42.0 44.5 48.6 49.1
Prec. (W/A) 6/6 4/4 6/6 4/4 6/6 4/4 6/6 4/4 6/6 4/4 6/6 4/4 6/6 4/4 6/6 4/4 6/6 4/4 6/6 4/4 6/6 4/4 6/6 4/4
MinMax 9.2 - 32.9 - 31.9 - 10.7 - 37.5 - 36.1 - 10.9 - 38.6 - 37.3 - 11.2 - 44.7 - 42.8 -
Percentile 10.9 - 33.5 - 32.0 - 12.0 - 38.0 - 36.3 - 12.3 - 39.0 - 37.5 - 14.0 - 45.4 - 43.1 -
OMSE 11.9 - 33.9 5.4 33.1 7.4 13.5 - 38.4 6.1 37.5 7.8 15.0 - 39.6 6.2 38.6 7.7 16.6 - 45.9 6.8 44.5 8.3
PTQ4SAM-S 15.4 - 35.7 18.1 36.0 24.1 17.4 - 40.0 20.6 40.3 26.7 17.9 - 41.0 20.9 41.3 27.3 20.4 - 47.7 23.1 48.1 30.5
SAQ-SAM 29.8 2.8 35.7 20.9 36.1 25.0 34.0 2.1 40.0 24.2 40.2 28.4 36.4 2.6 41.1 24.3 41.3 28.2 39.4 3.5 48.0 27.8 48.2 31.6
AdaRound 23.1 - 34.3 8.7 33.7 14.5 26.4 - 38.9 11.1 38.3 16.7 27.2 - 39.9 8.0 39.4 16.3 31.2 1.2 46.6 8.8 46.0 18.2
BRECQ 24.1 - 34.2 10.7 33.7 15.1 26.1 - 38.9 12.0 38.3 16.3 27.9 - 39.9 11.1 39.5 15.5 31.8 3.6 46.6 12.3 46.0 17.6
QDrop 29.3 13.0 35.2 22.6 36.3 32.3 33.6 13.3 39.7 25.3 40.4 35.8 34.3 13.2 40.5 25.8 41.4 36.5 38.9 11.2 47.5 27.5 48.3 41.7
PTQ4SAM-L 30.3 16.0 35.8 28.7 36.5 33.5 34.3 18.4 40.3 31.6 40.7 37.6 35.1 17.3 41.2 32.1 41.6 38.4 40.4 14.4 48.3 36.6 48.7 43.9
SAQ-SAM 32.0 27.7 35.9 34.5 36.6 35.9 35.9 30.3 40.3 39.0 40.9 39.9 37.1 31.8 41.3 39.8 41.8 40.7 42.4 33.8 48.3 46.3 48.9 47.4
Table 2: Results of oriented object detection on DOTA dataset. Our SAQ-SAM shows consistent superiority in both comparison groups.
SAM-B SAM-L SAM-H
Method FP W6A6 W4A4 FP W6A6 W4A4 FP W6A6 W4A4
Percentile 10.35 2.62 60.98 10.06 60.08 14.55
OMSE 18.89 1.11 61.55 23.50 61.11 17.49
PTQ4SAM-S 19.90 0.58 60.88 10.25 62.99 40.14
SAQ-SAM 63.93 51.84 4.63 64.35 62.70 27.83 64.49 63.35 41.63
AdaRound 34.05 - 63.44 23.18 62.73 24.45
BRECQ 34.40 - 63.60 26.89 62.58 25.98
QDrop 59.27 41.96 63.86 50.11 62.83 55.87
PTQ4SAM-L 60.33 44.18 63.91 56.29 64.36 56.01
SAQ-SAM 63.93 60.35 44.39 64.35 64.38 58.03 64.49 64.58 60.17
4 Experiments
@@ 4.1 Experimental Setup
Tasks and Datasets. Our experiments include three prevalent tasks with SAM: instance segmentation, oriented object detection, and semantic segmentation.
  • In instance segmentation, various object detectors provide prompting boxes for SAM to obtain segmentation masks, including Faster R-CNN (faster-rcnn), YOLOX (yolox), H-Deformable-DETR (detrs), and DINO (dino), with MS-COCO (coco) as evaluation dataset.
  • In oriented object detection, with a detector generating horizontal boxes as prompts, SAM predicts masks that are subsequently transformed into resulting rotated boxes. We do experiment on DOTA (dota) dataset with FCOS (tian2019fcos) detector.
  • In the semantic segmentation, SAM is employed to enhance the quality of segmentation masks produced by traditional semantic segmentor. We employ SegFormer (segformer) as the semantic branch and evaluate on ADE20K (ade20k) dataset.
@@
Implementation Details. We maintain consistant experimental setting with baselines (ptq4sam) for fair comparision. Specifically, the quantization scheme employs per-tensor asymmetric quantization for activations and per-channel asymmetric quantization for weights. The weights are calibrated using the MSE. The activation calibration set contains 32 images randomly sampled from the training dataset. For the PCC, we set threshold factor $\theta$ to 0.5. All QK activations (i.e., the inputs and outputs of Query Linear and Key Linear) are calibrated based on Attention Focus Overlap metric using the first sample. For PAR, the image encoder employs per-stage learning, while the mask decoder employs per-layer learning with 2000 iterations. Exceptionally, the additional final cross-attention block learns 10000 iterations due to its large loss. In general, compared to PTQ4SAM settings, which carry per-block learning with 20000 iterations, our method requires much lower time cost.
Baseline Methods. We take advanced PTQ4SAM as baseline, because both are dedicated to refining mask decoder quantization. For comprehensive comparison, we follow the PTQ4SAM settings and divide the methods into two groups according to whether there is a learning process. For statistic-based PTQ methods, we integrated PCC into PTQ4SAM-S, which is referred to as SAQ-SAM. And for learning-based PTQ, we build on PTQ4SAM-L and modified the reconstruction method from QDrop (qdrop) to our PAR, with PCC as pre-calibration process.
4.2 Performance Evaluation
@@
Instance Segmentation. In the task of instance segmentation, as shown in Table 1, our method consistently shows superior performance. Compared to baseline PTQ4SAM-S, SAQ-SAM achieves a nearly twofold improvement for 6-bit SAM-B, For SAM-L and SAM-H, the advantages are particularly evident in low-bit settings. For example, SAQ-SAM increases the mAP of 4-bit SAM-L from 23.1% to 27.8% with DINO. For reconstruction methods, compared to PTQ4SAM-L, our methods achieve significant improvement with lower computational costs. For example, SAQ-SAM improves the 4-bit SAM-B by 14.5% mAP and achieves near-lossless accuracy for 4-bit SAM-L and SAM-H.
Oriented Object Detection. In the oriented object detection task, as shown in Table 2, our method also yields excellent performance, demonstrating generality for accurate segmentation and orientation of small targets. For example, in statistic-based setting, our SAQ-SAM boost 6-bit SAM-B performance to a usable level (from 19.9% to 51.84%). In learning-based setting, SAQ-SAM improves the mAP of 4-bit SAM-L from 56.29% to 58.03%.
Table 3: Results of semantic segmentation on ADE20K dataset. Our SAQ-SAM demonstrate effective in improving segmentation performance.
SAM-B SAM-L
Method FP W6A6 W4A4 FP W6A6 W4A4
PTQ4SAM-S 31.16 31.08 33.48 22.64
SAQ-SAM 33.15 32.90 31.29 33.61 33.55 25.64
AdaRound 32.34 31.78 32.99 31.97
BRECQ 32.27 31.78 33.04 31.98
QDrop 32.57 31.79 33.58 32.67
PTQ4SAM-L 32.65 31.85 33.66 32.82
SAQ-SAM 33.15 33.04 32.53 33.61 33.63 33.30
Semantic Segmentation. In the task of semantic segmentation, the accuracy degradation caused by quantization is not significant due to the sophisticate post-processing steps. Nevertheless, as shown in Table 3, our methods also show superiority. For example, SAQ-SAM increases the accuracy of 6-bit SAM-B by 1.74% mIoU, even surpassing the learning-based PTQ4SAM-L. And SAQ-SAM further boosts the performance close to FP.
@@ 4.3 Ablation Study
With PTQ4SAM-L as baseline, we explore the effectiveness of SAQ-SAM components. As shown in Table 4, while PAR outperforms the baseline, when PCC act as pre-calibration for PAR, the combination of both achieves the best performance. See Appendix C.2 for more analysis.
Table 4: Ablation study of SAQ-SAM components.
Method SAM-B SAM-L SAM-H
Detector PAR PCC W6A6 W4A4 W6A6 W4A4 W6A6 W4A4
34.3 18.4 40.3 31.6 40.7 37.6
35.7 26.2 40.2 38.9 40.8 39.4
YOLOX 35.9 30.3 40.3 39.0 40.9 39.9
40.4 14.4 48.3 36.6 48.7 43.9
42.3 30.2 48.3 46.1 48.8 47.4
DINO 42.4 33.8 48.3 46.3 48.9 47.4
Attention Focus Threshold Setting of PCC. In the proposed PCC module, we use the threshold factor $\theta$ to define the salient region of attention, which is typically set at 0.5. To explore the impact of this hyperparameter setting, we perform a grid search on it. As shown in Figure 6 (a), with changes in $\theta$, the performance of PCC consistently outperforms the baseline method, albeit with slight fluctuations, where 0.5 is a modest choice. Notably, our method demonstrates exceptional robustness to the setting of $\theta$.
Figure 6: Ablation study of PCC and PAR. Figure (a) shows PCC is robust to the setting of $\theta$. Figure (b) shows per-stage PAR achieves optimal efficiency-accuracy trade-off.
Efficiency-Accuracy Study of PAR. To further verify the superiority of PAR, we explore its time consumption and performance with different granularity of optimization units, and compare it with QDrop-based reconstruction. As shown in Figure 6, per-stage scheme reduces the time cost of parameter learning, as well as boosting reconstruction performance, where per-stage PAR obtains optimal efficiency-accuracy trade-off. Furthermore, PAR performs better than QDrop under all granularity levels, highlighting the contribution of image-prompt interactions. The detailed experimental results can be found in the Appendix C.1.
PCC could serve as an orthogonal technique for combining with other advanced quantization methods. We present the improvement with RepQ-ViT and QDrop in the Appendix B.3. Since our method does not introduce any additional model components, the compression and acceleration benefits of quantized SAM are consistent with that reported in PTQ4SAM paper. See Appendix A.1 for details.